find the equation of a line passing through the points P[3,4] and Q[-5,8]

Grade:6

2 Answers

haritosh tripathy
6 years ago
find the slope in two point form that is=-1/2then the equation will be-----=(y-4)=-1/2(x-3)=2y+x-11=0 solved.
Punit
6 years ago
The slope of the points is=8-4/-5-3=-1/2. Now, y-4=-1/2(x-3), from here we will get the equation x+2y-11=0